You are the new marketing director for an up & coming Broadway theater. The theater just finished a renovation in order to more effectively compete with the completion. The renovation cost $75,000. It included an overall update of the theater along with additional seating options. It is your job to determine the average ticket prices for the 2020 season. Your biggest competitor is located closer to the theater district by 5 blocks. It is viewed by the consumer that theater closer to the district should be priced higher. Roughly $10 more per block. Their average ticket cost during the 2019 season was $100. However, the competing theater is not updated. You feel that the updated theater is worth $20 additional to each ticket. Determine the VTC for the 2020 average ticket cost.
Answer #1
Average ticket prices for the 2020 season :
Average ticket cost in 2019 100
Given,
an updated theater closer to the district should be priced higher by
per block 20
The difference in no.of blocks between this theater & the competitor is 5 blocks
Suppose the price/ticket of the competitor is x
Previously, before renovation , cost/tkt.of this theatre = (x-(5*10))= (x-50)
Now, after renovation , cost/ ticket(as the mgr. feels)= (x-50)+20
so, after renovation, the variable portion of the ticket (due to the location of the theatre, 5 blocks further away from the competitor)= -50+20=-30
ie. The variable portion (of costing a ticket)due to the the location has reduced by $ 20 to $ 30 per ticket--after renovation.
The answer is:
VTC for the 2020 average ticket cost = $ 30